Show that for s (semiperimeter), r(radius of inscribed circle), and R (radius of circumscribed circle), abc = 4srR.
Answer by venugopalramana(3286)

we got area of a triangle =(1/2)b*c* Sin A = r*s
we also have the formula that a/Sin A=2R..or...Sin A =a/2R..using this we get
(1/2)b*c* Sin A = (1/2)*b*c*a/2R=a*b*c/4R = r*s...or a*b*c=4R*r*s